OpenStudy (anonymous):

change 4.98 radians to degree measure?

OpenStudy (anonymous):

Ti change the mesure of an angle from the radians x to degrees y we have : \[\Large y=\frac{x\times 180}{\pi}\]

OpenStudy (anonymous):

\[\bf \pi \ radians = 180 \ degrees \implies 1 \ radian = \frac{180}{\pi} \ degrees\]Multiply both sides by 4.98.

OpenStudy (anonymous):

got it